OpenStudy (anonymous):

x^2-(a-3)x+a-2=0 has two integral solutions. List all possible 'a's.

OpenStudy (owlcoffee):

are you supposed to use integrals or just solve that equation?

OpenStudy (kc_kennylau):

If b^2-4ac < 0, no solutions exist If b^2-4ac = 0, one solution exists. If b^2-4ac > 0, two solutions exist.

OpenStudy (owlcoffee):

okay so in other words, we have to study the behaviour of "a". So let's write it on those terms: \[-(a-3)^{2}-4(a-2)\] let's expand that binome and apply distributive on the other hand: \[-a ^{2}+6a-9-4a+8\] let's operate similar terms: \[-a ^{2}+2a-1\] now I want to know what value of "a" makes that 0 so: \[-a ^{2}+2a-1=0\] and we get that a=1. so we found a value of a that determines a horizon in the discriminant part, so we can say: when a<1 \[-(a-3)^{2}-4(a-2) <0 \] there are no solutions for the equation when a=1 \[-(a-3)^{2}-4(a-2) = 0\] there is only one solution for the equation when a>1 \[-(a-3)^{2}-4(a-2)>0\] there are two solutions for the equation. we the list of all the "a" values that give us 2 solutions are [1,+inf)

OpenStudy (anonymous):

*

OpenStudy (owlcoffee):

oooh, why u stop? :( That was very interesting.

OpenStudy (anonymous):

we want t come up with 2 integral solutions, so discriminant of quadratic must be a positive square, in mathematical words\[\Delta=(a-3)^2-4(a-2)=a^2-10a+17=m^2\]

OpenStudy (anonymous):

\(m\) is a positive integer (why?) and\[a^2-10a+17=(a-5)^2-8=m^2\]\[a=5\pm\sqrt{m^2+8}\]this is necessary but not sufficient, let's check the solutions\[x_1,x_2=\frac{a-3\pm\sqrt{a^2-10a+17}}{2}=\frac{2\pm\sqrt{m^2+8}\pm m}{2}\]

OpenStudy (anonymous):

now \(\pm m\pm\sqrt{m^2+8}\) must be an even integer (why?) it's possible only if \(m=1\) (why?)

OpenStudy (anonymous):

finally \(a=8,2\) :)

OpenStudy (owlcoffee):

hohohoho, I'll take note of that method, looks interesting to do some profound research.